# A review and experiences from quantum physics educators

### v1

@@ Abstract
Students often enter physics classrooms with deeply ingrained misconceptions, typically stemming from common intuition and everyday experiences. These misconceptions present significant challenges for educators, as students are often resistant to information that conflicts with their preconceptions. As a result, traditional instructional methods often fail to address misconceptions. The first aim of this manuscript is to summarize the existing literature on misconceptions in university physics. This resource for instructors reviews misconceptions’ sources, diagnoses, and remediation strategies. Like most physics education research, the majority of this literature has concentrated on classical physics. However, quantum physics poses unique challenges because its concepts are far removed from everyday experiences and intuition. This uniqueness signals the need to ask how well the strategies developed for addressing misconceptions in classical physics apply to quantum physics. This need is underscored by the recent surge of people from diverse backgrounds entering quantum physics because of the growing significance of quantum technologies in fields such as computing, cryptography, and materials science. To help answer this question, we conducted in-depth interviews with quantum physics instructors at the University of Waterloo who have collectively taught over 100 university quantum physics courses. These interviews explored the nature of common misconceptions in quantum physics, their origins, and effective instructional techniques to address them. We highlight specific misconceptions, such as misunderstanding of entanglement and spin, and successful teaching strategies, including “misconception-trap quizzes.” We integrate insights from the literature review with our interview data to provide an overview of current best practices in addressing physics misconceptions. Furthermore, we identify key research questions that warrant further exploration, such as the efficacy of multi-tier tests in quantum physics and developing a cohesive quantum curriculum. This paper aims to inform educators and curriculum developers, offering practical recommendations and setting a research agenda to improve conceptual understanding in classical and quantum physics.
@@ I Introduction
Physics is filled with concepts that challenge common intuition and are difficult for students to grasp. For example, many students mistakenly believe that heavier objects fall faster than lighter ones or that a constant force is needed to keep an object in motion. Such misconceptions often arise from everyday experiences and intuitive reasoning, leading to persistent barriers in learning even after formal instruction [1, 2].
Misconceptions can be more detrimental than gaps in knowledge. This is because students tend to resist new information that conflicts with their existing beliefs, resulting in cognitive dissonance and frustration [3, 4]. Therefore, effective teaching strategies that identify and rectify these misconceptions are essential for fostering a conducive learning environment.
However, physics educators face significant challenges in identifying and rectifying these misconceptions. Traditional instructional methods often fail to address the underlying intuitive beliefs that give rise to misconceptions [5, 6]. Not all traditional teaching methods teach for conceptual change [7]. Consequently, educators must employ innovative and targeted strategies to help students develop a more accurate conceptual framework. Fortunately, a sizeable body of research has been conducted on evidence-based strategies for identifying and addressing misconceptions. The first aim of this manuscript is to review the existing literature on misconceptions in university physics, focusing on their sources, diagnosis, and remediation strategies.
Most physics education is centred on classical physics; naturally, most physics education research also focuses on classical physics. This trend extends to the study of misconceptions. For instance, Gilbert and Watts [1] reviewed 33 studies from 1972 to 1982 on identifying misconceptions, and Soeharto et al.  [2] reviewed another 111 articles from 2015 to 2019. Of the over 40 misconceptions identified in these studies, none were on quantum physics, highlighting a potential gap in the research.
Quantum physics introduces a unique set of challenges due to its abstract and non-intuitive nature [8, 9, 10]. Phenomena such as wave-particle duality and entanglement are far removed from everyday experiences and defy classical logic. This uniqueness underscores the need for increased attention to identifying and remedying misconceptions in quantum physics. The importance of addressing these misconceptions is further amplified by the growing significance of quantum technologies in fields such as computing, cryptography, and materials science. This growth has resulted in learners from a broad range of academic backgrounds looking to learn the basics of quantum physics.
Strategies designed to diagnose and remedy misconceptions in classical physics often rely on analogies and experiences from the macroscopic world, which do not always translate well to quantum phenomena. Consequently, while it is essential to determine which ideas and tools from classical physics education can be adapted for quantum physics, it is equally important to recognize that some may not directly apply. To explore the applicability of these strategies, we conducted interviews with quantum physics instructors at the University of Waterloo, a leading institute for quantum physics which houses the Institute for Quantum Computing and has strong affiliations with the Perimeter Institute for Theoretical Physics. Collectively, these instructors have taught over 100 quantum physics courses. Since they are on the front lines of teaching and observing student misconceptions, they provide valuable insights into the effectiveness of existing educational tools and the unique challenges faced in teaching quantum concepts. These interviews aim to identify which educational strategies can be ported over from classical to quantum physics. This is the second aim of this manuscript.
The rest of the manuscript is organized as follows. Section II presents our review of the literature on addressing misconceptions in university physics. Instead of providing an exhaustive survey, we highlight key results and illustrative examples of each. A summary of the interviews with quantum physics instructors is presented in Section III. By integrating insights from these interviews with the existing literature on classical physics misconceptions, we aim to provide an overview of best practices for addressing misconceptions in classical and quantum physics and identify research opportunities. This integrated discussion is presented in Section IV.
II Review
This section reviews the literature on misconceptions in university physics education, beginning with a review of the broader context of potential sources of misunderstanding (Section II.1). In Section II.2, we explore the origins of these misconceptions. Next, we examine methods for diagnosing misconceptions in Section II.3. Finally, in Section II.4, we discuss strategies for remedying these misconceptions. We propose that these remediation strategies can be categorized into three main types: experiment-based, discussion-based, and reflection-based. This categorization helps to identify common elements and key factors contributing to their effectiveness.
II.1 Misconceptions among other sources of misunderstanding
@@
The study of misconceptions begins with defining them precisely and distinguishing them from other sources of misunderstanding. In their 1983 review, Gilbert and Watts noted the initial lack of consensus on the definition of a misconception within the field [1]. We review two well-known typologies for categorizing misunderstandings and simplify them into a two-bin system: misconceptions, stemming from incorrect existing knowledge, and gaps in prior knowledge.
@@
Our category “gaps in knowledge” aligns with 2a, and “misconceptions” with 1a-1d and 2b-2c. Which binning 1e or 2d goes under may depend on context, but it seems more likely to be a knowledge gap. For example, translating a word incorrectly is a vernacular misunderstanding that can likely be corrected easily without leading to deeply ingrained incorrect beliefs. Having clearly defined misconceptions and tied our definition to the existing ones used in the literature, we turn to their sources.
II.2 Sources of misconceptions
Misconceptions can arise from three sources:
@@ II.2 Sources of misconceptions
II.3 Diagnostic Methods
Being mindful of sources of misconceptions will reduce the number of students’ misconceptions, but some will emerge. To help students address these misconceptions, we need tools to diagnose them. The commonly used tools are interviews, open-ended questions, certainty of response index (CRI) tests, and multiple-tier tests. The first two strategies are typically used in courses with few students, and the latter two in courses with many students.
Interviews and open-ended questions
@@ II.3 Diagnostic Methods
CRI and multiple-tier tests
@@
II.4 Remediation Strategies
After a misconception is identified, it can be addressed. This requires effective strategies grounded in educational research to foster conceptual change [7]. In this section, we propose a typology that categorizes these strategies into three groups: reflection-based, discussion-based, and experiment-based approaches. These categories highlight the underlying principles shared amongst various strategies. Examples for each category are summarized in Table 1.
@@ II.4 Remediation Strategies
Reflection-based strategies
@@
Discussion-based strategies
@@
Experiment-based strategies
@@
III Experiences from quantum physics educators
Most of the tools for identifying and addressing misconceptions were developed for classical physics. Since quantum physics education presents unique challenges, we need to consider the applicability of these tools to this subject. We conduct such an evaluation through long-form interviews with quantum physics instructors at the University of Waterloo. This study aims to identify misconceptions in quantum physics and assess effective tools for diagnosing and rectifying these misconceptions.
III.1 Methodology Overview
First, we outline our methodology for identifying instructors and conducting interviews. The University of Waterloo, renowned for its quantum physics research and extensive course offerings, was chosen for this study. It houses the Institute for Quantum Computing and is closely affiliated with the nearby Perimeter Institute for Theoretical Physics. A testament to the university’s focus on quantum physics is the volume of courses it offers on the subject. Listed in Appendix A, the university offers 16 undergraduate and 37 graduate courses on solely quantum physics and 9 undergraduate courses covering quantum physics.
We identified faculty members who taught the 16 undergraduate courses via online course listings and departmental contacts. Emails were sent to these faculty members requesting interviews, resulting in nine agreements. Through these nine, another faculty member was identified to give an interview about their experience teaching one of the nine undergraduate courses partially covering quantum physics. Additional interviews were secured with two individuals from the Institute for Quantum Computing and Perimeter Institute with years of experience teaching quantum physics to undergraduate students. The twelve instructors have collectively taught over 100 university quantum physics courses and many summer schools, reaching thousands of students.
With no detailed guide for conducting physics education research interviews, we relied on guides from other fields [43, 44, 45, 46]. The interviews aimed to identify misconceptions in quantum physics and effective tools for addressing them. Five core questions were developed, with flexibility for follow-up questions based on responses. Verbal consent for recording was obtained, and confidentiality and withdrawal rights were explained to each interviewee. Interviews were scheduled for 30 minutes each, conducted via Zoom, and many were extended beyond the scheduled time. The insights gathered from these interviews are synthesized and presented in a cohesive narrative in the following subsection.
III.2 Synthesis of Insights
We present each question and list the various answers instructors gave. With few exceptions, each answer is an amalgamation of comments given by multiple instructors.
@@ III.2 Synthesis of Insights
What are the sources of these misconceptions?
@@
What are your strategies for identifying misconceptions?
@@
What are your strategies for addressing misconceptions?
@@
Other miscellanenous comments
@@
Specialization in quantum physics:.—The importance of tailoring quantum mechanics education to the audience and its goals was emphasized. Mathematics students may approach quantum mechanics primarily through the lens of linear algebra with additional rules. These students could go on to prove useful results about, for example, quantum algorithms without needing knowledge of, say, scattering. Quantum physics research requires a diverse set of skills, and specialization is necessary. Whether this level of specialization is relevant at the undergraduate level is unclear. However, the question is then raised on what to include in undergraduate curriculums as the number of quantum physics education courses expands.
III.3 Limitations of the study
In this work, we focused on one institute that is a hot spot for quantum physics education. This intense concentration of experiences made Waterloo a prime location for analyzing quantum physics education. However, Waterloo is one of Canada’s most competitive physics programs and is recognized as a global leader in quantum physics. As such, the student population is likely biased to have a stronger physics background than the average student. Furthermore, the instructors likely understand the content more deeply than those at a randomly chosen university. Future research should aim to include a larger and more diverse sample of instructors from various institutions and cultural backgrounds to enhance the generalizability of the findings.
The study relies heavily on qualitative data from interviews, which, while rich in detail, may introduce subjectivity and bias. The interpretations and conclusions drawn are influenced by the educators’ and interviewer’s perspectives and experiences. Including quantitative measures, such as surveys or pre- and post-assessments of student performance, could provide a more balanced view and validate the effectiveness of the suggested strategies.
IV Discussion
Physics educators often face challenges due to ingrained student misconceptions. Classical and quantum physics present unique hurdles, as highlighted by the literature and reinforced by our interviews with quantum physics educators at the University of Waterloo. This final section synthesizes many of the conceptual insights from our review and discussions and identifies key research questions. However, one should refer to the main text for a detailed list of strategies.
IV.1 Integration of insights from literature and interviews

### v2

# Experiences from quantum physics educators

@@ Abstract
Students often begin physics courses with misconceptions rooted in everyday experience and intuition, which can be resistant to change. While research has identified strategies for addressing misconceptions across physics, it remains unclear whether different domains, like classical and quantum physics, require different approaches. On one hand, quantum concepts are further removed from daily experience, possibly requiring specialized strategies. On the other, all physics must be learned rather than innately understood, and classical physics already contains many counterintuitive ideas, suggesting the same strategies may be equally valid. To explore this question, we first develop a structure to organize the existing literature on addressing misconceptions in physics education. We identify 126 studies, which we group into six categories, three of which include further subcategories. Rather than offering a comprehensive literature review, this scheme provides a structural lens through which to compare how various strategies align. We then take an instructor-centered approach, using our framework to guide interviews with quantum physics instructors. We interview 12 instructors from the University of Waterloo’s Institute for Quantum Computing and the Perimeter Institute who have collectively taught over 100 quantum courses. Our findings suggest that quantum physics instructors find strategies similar to those used in classical physics effective for identifying and addressing misconceptions. This highlights the potential to adapt, rather than replace, existing instructional tools to better meet the conceptual challenges of quantum physics. Our work provides practical insights for educators and curriculum designers aiming to support deeper understanding in quantum mechanics.
@@ I Introduction
Physics education research has long grappled with the challenge of student misconceptions (also known as “alternative conceptions,” “conceptual misunderstandings,” or “preconceptions”). These are deeply held, often intuitive beliefs that conflict with formal physics understanding. Rather than random errors, they represent robust mental models grounded in everyday experience and persist despite formal instruction. Over the years, researchers have identified many of these common misconceptions and developed evidence-based strategies to confront and revise them [1, 2, 3].
An open question is whether classical and quantum physics courses demand distinct strategies for identifying and addressing misconceptions. On one hand, quantum concepts are more abstract and less tied to everyday experience, potentially requiring novel instructional approaches. On the other hand, all physics—classical or otherwise—must be learned rather than innately grasped, and classical physics also contains counterintuitive ideas, implying that existing methods may still be effective. While substantial work has identified misconceptions in both classical and quantum domains, few studies have systematically compared how these misconceptions emerge and are addressed across fields, leaving instructors without clear guidance on whether quantum physics instruction truly demands unique strategies or merely different emphases.
This study helps address that gap by investigating how highly experienced quantum physics instructors view the strategies for dealing with misconceptions, and whether these strategies must differ from those commonly used in classical contexts. First, we construct a framework to organize the existing literature on misconception-oriented instruction, drawing on 126 studies across various subfields. This framework then guides a qualitative investigation of expert teaching practice: we conduct structured interviews with 12 quantum physics instructors at the University of Waterloo’s Institute for Quantum Computing and the Perimeter Institute, who collectively have taught over 100 courses.
Our research is guided by three main questions:
  • RQ1:
    What sources and strategies have been identified in the physics education literature for diagnosing and addressing student misconceptions, and how can they be systematically categorized?
  • RQ2:
    Do instructors find these strategies effective in quantum physics courses?
  • RQ3:
    How do the sources and nature of misconceptions in quantum physics compare to those in classical physics from the perspective of instructors?
By addressing these questions, we aim to clarify whether teaching quantum physics requires fundamentally new pedagogical methods, or whether existing approaches can be adapted to meet its unique conceptual demands.
To situate our study, we first highlight key contributions in the literature on student misconceptions in quantum physics instruction. A foundational review by Singh and Marshman (2015)[4] synthesized research on common difficulties in upper-level quantum mechanics, emphasizing prevalent reasoning errors, conceptual overgeneralizations, and the importance of research-driven instructional strategies. Complementing this, a 2017 review by Krijtenburg-Lewerissa et al.[5] examined teaching practices at the secondary and introductory undergraduate levels, identifying shared misconceptions, gaps in pedagogical approaches, and a lack of validated assessment tools.
Since these reviews, the field has seen a marked increase in targeted studies. Several have focused on specific conceptual challenges students face, such as difficulties with separation of variables [6], solving bound and scattering state problems [7], and applying delta functions in quantum mechanics [8]. Parallel work has explored strategies for conceptual remediation across a range of foundational topics, including single-photon interference in Mach-Zehnder interferometers [9], the formalism and postulates of quantum mechanics [10], the interpretation of the double-slit experiment [11], and degenerate perturbation theory in the context of the Zeeman effect [12]. These efforts have contributed to the development of research-based tools such as inquiry-driven tutorials, interactive simulations, and conceptual surveys designed to enhance learning and measure instructional effectiveness. In addition, some studies have adopted epistemological frameworks to analyze how students frame and engage with quantum mechanics problems [13, 14].
Our study expands on this body of research in two key ways. First, we center our analysis on the perspectives of experienced instructors, offering a complementary viewpoint. Second, we compare strategies across classical and quantum domains to clarify the extent to which established strategies remain effective when addressing quantum physics.
The rest of this paper is organized as follows: Section II explains the methodology for identifying relevant literature and conducting instructor interviews. In Sec.III, we clarify the distinction between misconceptions and other sources of misunderstanding. SectionIV details our classification scheme for categorizing the sources of misconceptions, diagnosis methods, and remediation strategies, with illustrative examples. Section V then presents the interview findings, highlighting instructors’ insights into the nature of quantum misconceptions and the perceived efficacy of various instructional strategies. Finally, Section VI synthesizes these findings to evaluate the transferability of classical teaching approaches to quantum contexts and outlines directions for future research.
II Method
II.1 Literature Search
To conduct a systematic review of misconceptions in university physics education, we employed a multi-stage literature search. Our approach combined database searches, structured screening, and citation-based expansion. First, we search the Web of Science database using the exact string:
ALL = ("misconception" OR "alternative conception" OR "preconception" OR "conceptual misunderstanding") AND ALL = ("physics")
This generated $419$ results and was carried out in January 2024. The first exclusion criteria was
  • The article had not been published in a peer-reviewed journal.
This reduced the list to $219$ entries. Each of these $219$ was then independently screened based on the following exclusion criteria.
  • The article is not in the field of physics education research.
  • The article is unrelated to addressing misconceptions.
These criteria reduced the list to $102$ entries. To identify overlooked studies, we employed forward and backward citation tracking:
  • Backward snowballing: Reviewing references of selected studies.
  • Forward snowballing: Using Google Scholar’s citation function to find newer papers citing the selected studies.
This added $24$ papers, bringing the total to $126$. Each of these $126$ papers was then sorted into one of four categories: Sources of misconceptions, Diagnostic methods, Remediation strategies, and Other. The categories were further subdivided into subcategories (see Fig. 1). Each paper, with each subcategory and a brief summary is provided in Appendix B.
Figure 1: Overview of our organizational scheme for the literature on misconceptions in physics. Numbers indicate the count of papers in each category. For brief summaries of each paper, see Tables 2 to 7 in Appendix B.
II.2 Interviews
II.2.1 Participant Selection
We focused on the University of Waterloo due to its recognized strengths in quantum physics education and its affiliation with the Perimeter Institute. The university offers 16 undergraduate and 37 graduate courses on solely quantum physics and 9 undergraduate courses covering quantum physics (listed in Appendix A). Initial participants were identified through official course listings; they were contacted by email using a standardized invitation. We also used snowball sampling by asking interviewees to recommend colleagues with relevant experience.
In total, 12 instructors (nine current faculty members and three experienced guest lecturers for summer schools) agreed to participate. Each participant had taught quantum physics in a university setting for a minimum of five years, ensuring substantial teaching experience.
II.2.2 Interview Protocol
We designed our interview protocol following guidelines from qualitative research in education [15, 16, 17, 18]. The final protocol included five core, open-ended questions addressing:
  • What are the sources of these misconceptions?
  • What are your strategies for identifying misconceptions?
  • What are your strategies for addressing misconceptions?
We asked follow-up questions as relevant, while ensuring all core questions were covered to maintain consistency across interviews. Each interview was scheduled for 30 minutes via Zoom, though many extended up to 45 minutes or more based on participant engagement.
Interviews were audio-recorded and professionally transcribed. All participants provided recorded verbal consent. We assured them of confidentiality; hence, all identifying information was removed from transcripts.
II.2.3 Strengths and Limitations of the Interview Approach
Interviews with lecturers are not the standard method for identifying students’ difficulties. As a non-standard approach, we carefully examine its strengths and weaknesses. One strength is that lecturers, through years of teaching, develop an intuitive understanding of common student struggles, providing a broad overview of recurring challenges. Additionally, unlike standardized tests or surveys, interviews allow lecturers to reflect on not only errors in assessment but also patterns in classroom engagement, misconceptions, and areas where students frequently seek help. Moreover, interviews can be more time-efficient than large-scale student surveys or detailed diagnostic assessments. However, this method has notable weaknesses. Lecturer recollections of student difficulties may be influenced by selective memory or biases. The identification of challenges is subjective and may vary between lecturers based on their personal experiences, teaching styles, and interactions with students. Furthermore, some student difficulties may not be apparent to lecturers, particularly if students do not actively express them in class. Lastly, because this method does not directly involve students, it risks overlooking nuanced challenges that students themselves perceive but lecturers do not recognize.
To ensure the validity of our findings with this strategy, we implemented several procedures. First, we triangulated interview data with a literature review to strengthen our conclusions. Second, we standardized the interview process by using structured questions and a systematic approach, focusing on specific topics and common misconceptions to reduce arbitrariness. Third, we gathered insights from multiple lecturers teaching the same subject to identify consistent patterns and enhance the robustness of our findings. These measures help mitigate the limitations of lecturer interviews and improve the reliability of our conclusions.
III Defining “Misconception”
@@
Before presenting our main results, we define what is meant by a “misconception” by comparing it against other sources of misunderstanding. In their 1983 review, Gilbert and Watts noted the initial lack of consensus on the definition of a misconception within the field [1]. We review two well-known typologies for categorizing misunderstandings and simplify them into a two-bin system: misconceptions, stemming from incorrect existing knowledge, and gaps in prior knowledge.
@@
Our category “gaps in knowledge” aligns with 2a, and “misconceptions” with 1a-1d and 2b-2c. Which binning 1e or 2d goes under may depend on context, but it seems more likely to be a knowledge gap. For example, translating a word incorrectly is a vernacular misunderstanding that can likely be corrected easily without leading to deeply ingrained incorrect beliefs.
IV Categorizing the literature
This section addresses RQ1, by creating broad categories to organize the literature based on the papers we found via our literature search. This categorization helps to identify common elements and key factors contributing to their effectiveness. In Section IV.1, we explore the origins of these misconceptions. Next, we examine methods for diagnosing misconceptions in Section IV.2. Finally, in Section IV.3, we discuss strategies for remedying these misconceptions. For each subcategory we present at least one illustrative example.
IV.1 Sources of misconceptions
The literature we found on sources of misconceptions is presented and sorted in Table 2. Misconceptions can arise from three sources:
@@ IV.1 Sources of misconceptions
IV.2 Diagnostic methods
Being mindful of sources of misconceptions will reduce the number of students’ misconceptions, but some will emerge. To help students address these misconceptions, we need tools to diagnose them. The literature we identified on diagnostic methods is presented and sorted in Table 3. The commonly used tools are interviews, open-ended questions, certainty of response index (CRI) tests, and multiple-tier tests. The first two strategies are typically used in courses with few students, and the latter two in courses with many students.
IV.2.1 Interviews and open-ended questions
@@
IV.2.2 CRI and multiple-tier tests
@@
IV.3 Remediation strategies
After a misconception is identified, it can be addressed. This requires effective strategies grounded in educational research to foster conceptual change [2]. In this section, we propose a typology that categorizes these strategies into three groups: reflection-based, discussion-based, and experiment-based approaches. These categories highlight the underlying principles shared amongst various strategies. Examples for each category are summarized in Table 1. All of the literature we examined in this category is presented in Table 4 and 5.
@@ IV.3 Remediation strategies
IV.3.1 Reflection-based strategies
@@
IV.3.2 Discussion-based strategies
@@
IV.3.3 Experiment-based strategies
@@
V Interviews with quantum physics educators
Most of the tools for identifying and addressing misconceptions were developed for classical physics. Since quantum physics education may present unique challenges, we need to consider the applicability of these tools to this subject. We conduct such an evaluation through long-form interviews with quantum physics instructors at the University of Waterloo. These interviews aim to address RQ2 and RQ3. We present each question and list the various answers instructors gave. With few exceptions, each answer is an amalgamation of comments given by multiple instructors.
V.1 What are university students’ common misconceptions about quantum physics?
@@
V.2 What are the sources of these misconceptions?
@@
V.3 What are your strategies for identifying misconceptions?
@@
V.4 What are your strategies for addressing misconceptions?
@@
V.5 Besides misconceptions and gaps in knowledge, what else causes students to make mistakes?
@@
V.6 Other miscellanenous comments
@@
Specialization in quantum physics.—The importance of tailoring quantum mechanics education to the audience and its goals was emphasized. Mathematics students may approach quantum mechanics primarily through the lens of linear algebra with additional rules. These students could go on to prove useful results about, for example, quantum algorithms without needing knowledge of, say, scattering. Quantum physics research requires a diverse set of skills, and specialization is necessary. Whether this level of specialization is relevant at the undergraduate level is unclear. However, the question is then raised on what to include in undergraduate curriculums as the number of quantum physics education courses expands.
VI Discussion
Physics educators often face challenges due to ingrained student misconceptions. Whether classical and quantum physics requires unique strategies for identifying and addressing them is an open question. From the interviews we conducted, instructors seem to find the same strategies effective for quantum physics instruction as instruction of other courses. We conclude here with a brief summary of the insights this work gleaned from the literature search and interview and identifying open research questions.
VI.1 Integration of insights from literature and interviews